yeetantor:

Dan mixed 2 parts of milk and 3 parts of syrup to make a drink. How many parts of milk did Dan mix with each part of syrup?

jimthompson5910:

The question seems a bit vague. But I think they're asking how many parts of syrup goes with 1 part milk, and how many parts of milk go with 1 part syrup. Though I could be wrong.

yeetantor:

well idk really i dont know math i need your helppppp

yeetantor:

im so confused about this question

jimthompson5910:

If my assumption is correct, then think of it like an equation like so 2 parts milk = 3 parts syrup

jimthompson5910:

If we divide both sides by 2, then "2 parts milk" turns into "1 part milk". Agreed?

yeetantor:

agreed

jimthompson5910:

What happens when we divide both sides by 2

jimthompson5910:

So you'll need to divide 3 over 2

yeetantor:

ok

yeetantor:

1.5

jimthompson5910:

yes, so 1 part milk corresponds to 1.5 parts syrup 1.5 = 1 & 1/2

jimthompson5910:

So you can think of it like 1 cup milk goes with 1 and a half cups of syrup maybe

yeetantor:

so the answer is 1.5

jimthompson5910:

Oh wait I just reread the problem. I know what your teacher wants now

yeetantor:

yea

jimthompson5910:

" How many parts of milk did Dan mix with each part of syrup?" So think of it as "how many parts of milk go with 1 part syrup"

jimthompson5910:

again we have 2 parts milk = 3 parts syrup we'll divide both sides by 3 to turn the "3 parts syrup" into "1 part syrup"

yeetantor:

ok

yeetantor:

is that it

jimthompson5910:

what do you get

yeetantor:

what im i dividing

jimthompson5910:

2 over 3

yeetantor:

oh ok

yeetantor:

what is that the answer

jimthompson5910:

you tell me, what does your calculator say?

jimthompson5910:

2 parts milk = 3 parts syrup 2/3 parts milk = 3/3 parts syrup ?? parts milk = 1 part syrup What replaces the question marks

yeetantor:

6

jimthompson5910:

you don't multiply them

yeetantor:

oh hold on

yeetantor:

it says 0.666666

jimthompson5910:

yes

yeetantor:

im confused

jimthompson5910:

so roughly 0.667 parts milk correspond to 1 part syrup

yeetantor:

ok so what im still confused here are the question answers two over five two over three three over five three over two

jimthompson5910:

oh so you don't convert to decimal form you just keep it as a fraction

yeetantor:

yea

jimthompson5910:

so hopefully you see the answer at this point

yeetantor:

i dont

yeetantor:

2 over 3

jimthompson5910:

2 parts milk = 3 parts syrup 2/3 parts milk = 1 part syrup

jimthompson5910:

yes

yeetantor:

ok